Explain how credibility might affect the cost of recording inflation.
1
Expert's answer
2014-11-19T13:11:34-0500
The Lucas critique states that it is improbable to assume that wage setters would not consider changes in policy when forming their expectation. If the wage setters believes that the policymakers would be committed in decreasing the inflation rate, they would lower their expectations of inflation and this will lead to the decline in the rate of actual inflation without the need of prolonged recession. This can be explained with the help of the above-mentioned equation in which expected inflation is taken on the right: Өt=Өte-ἀ(ut-un). If the wage-setters look at the previous year's inflation rate and form their expectations accordingly, then inflation rate can be reduced only by accepting a higher rate of unemployment for some period. If Өte=Өt-1, from Өt-Өt-1=-ἀ(ut-un. Thus in order to achieve: Өt < Өt-1, it must be that ut > un) But if the wage-setters convince themselves that the rate of inflation will fall in the future from 9% to 5%, that is to say it was indeed going to be lower than the past, thus forming their expectations accordingly, then inflation would fall to 5% even if unemployment remains at natural rate of unemployment.
One of the most important constituents of successful disinflation is the credibility of monetary policy according to Thomas J. Sargent. It states that the beliefs of wage setters are affected if they feel that the central bank are religiously committed in reducing the rate of inflation. The way the wage-setters formed their expectations can only be changed with the help of credibility. The credibility view is that fast disinflation is likely to be more credible than slow disinflation. Credibility decreases the unemployment cost of disinflation. Therefore, the central bank should go for fast disinflation.